Fermi National Laboratory
 

Accelerator Update for May 27, 2009 - May 29, 2009

 

Wednesday May 27

The day shift began with the Tevatron (TeV) quench checkout continuing, with Operations monitoring the Recycler stash (192.1E10), with the Antiproton Source stacking (27.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

The D0 quench protection monitoring system investigation continued until 10:30 AM and was followed by studies aimed at understanding the quench. Near the end of the day shift another quench protection system problem, this time at TeV sector C2, halted studies from 3:30 PM into the evening shift.

The TeV sector C2 quench protection system problem was resolved by 5:30 PM and the studies continued until 6 PM.

Operations established beam to a new MTest experiment, T979, at 4:45 PM. They disabled beam at 6 PM.

Operations put the TeV into HEP shot setup and established store 7089 at 7:45 PM with an average initial luminosity of 242.2E30.

A recurring problem with a Linac extraction power supply caused downtime for all users from 9:45 PM until 11 PM.

Thursday May 28

The midnight shift began with Operations monitoring store 7089 (125.71E30) and the stash (247.8E10), with Pbar stacking (17.0E10), and with NuMI and MiniBooNE taking beam.

Operations established beam to MTest 6 AM.

Operations terminated store 7089 at 8:12 AM.

Operations put the TeV into HEP shot set-up by 8:23 AM.

Operations established store 7090 at 9:56 AM with an initial luminosity of 263.54E30.

Operations disabled beam to MTest at 6:30 PM.

Operations terminated store 7090 at 7:42 PM.

Operations put the TeV into HEP shot set-up by 7:54 PM.

Operations established store 7091 at 9:20 PM with an initial luminosity of 291.55E30.

Friday May 29

The midnight shift began with Operations monitoring store 7091 and the stash (120.5E10), with Pbar stacking (12.5E10), and with NuMI and MiniBooNE taking beam.

The Tevatron aborted and Operations received a Tevatron BBM alarm at 12:06 AM. The Master Beam Switch was taken and experts were called to investigate.

Beam was returned to NuMI, MiniBooNE and PBar at 12:59 AM.

D0 started a controlled access at 1:36 AM and completed it by 2:59 AM.

At 5:20 AM, experts determined that the cause of the Tev abort and BBM alarm was caused by a bad Fiber Abort Repeater at C1. The module was replaced and Tev studies have begun.

Operations established beam to MTest at 7:10 AM.

The Plans for Friday and the Weekend

The plan for today and for the weekend is to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/01/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory